OpenStudy (anonymous):

A compressed-air cylinder stands 93.0 cm tall and has internal diameter 18.2 cm. At room temperature, its pressure is 167 atm. How many moles of air are in the cylinder? And what volume would this air occupy at room temperature and 5 atm pressure? So I have done this problem, but keep getting it wrong. Any help would be great.

OpenStudy (anonymous):

I started out by finding the volume => V=(pi)(r)2(h) and then I plugged in everything in the equation PV=nRT => (167atm)(V)=n(.08206)(298K), but i'm not getting the right answer. Any idea where it went wrong?

OpenStudy (anonymous):

(167atm)(24.2L)=n*(0.08206 L*atm/mol*K)*(298) therefore n= 165.266 now using this n solve for the new volume.

OpenStudy (anonymous):

is that the wrong answer you got?

OpenStudy (anonymous):

obv. units for n= 165.266 moles

OpenStudy (anonymous):

Oh I see what I did wrong, I didn't change the units and instead of getting the correct answer 165 I kept getting 0.165. Thanks so much :)

OpenStudy (anonymous):

lol no problem!